# A Survey of Multimodal Large Language Model from A Data-centric Perspective

### v1

@@ 3.2.2. Deduplications
Existing deduplication methods consider exact duplication, approximate duplication and semantic duplication in sentence-level (or sequence-level) and document-level. Exact deduplication is the most simple way to remove duplication. For sentence or sequence level, exact deduplication consider to remove exact string matching between sentence (or sequence) (Raffel et al., 2020; Suárez et al., 2019). To improve computational and memory efficiency, Suffix Arrays (Manber and Myers, 1993; Lee et al., 2022) and Bloom Filters (Computer, 2023; Soldaini et al., 2024) are used to achieve parallelized linear time complexity. For document level exact deduplication, URL deduplication is considered to remove exactly same web pages (Penedo et al., 2023). Approximate deduplication methods mainly focus on document-level duplicates. Usually, Locality Sensitive Hashing (LSH)-based MinHash (Broder, 1997) or SimHash (Charikar, 2002) methods are used to remove approximately duplicate documents (Computer, 2023; Lee et al., 2022; Penedo et al., 2023). These methods can achieve document-level deduplication with linear time and space complexity of the document number and can be implemented in a highly distributed setting (Lee et al., 2022). Apart from hashing-based approximate deduplication, recently there are some work consider to leverage pre-trained foundation models as semantic embedding metric for document-level embedding (Silcock et al., 2022; Kaddour, 2023; Abbas et al., 2023; Tirumala et al., 2023). These Semantic deduplication methods consider using Sentence-BERT (Reimers and Gurevych, 2019) MPNET (Song et al., 2020) for embedding (Silcock et al., 2022), E5-Large (Wang et al., 2022b) for embedding (Kaddour, 2023), and OPT-125M (Zhang et al., 2022) for embedding (Abbas et al., 2023; Tirumala et al., 2023). These semantic deduplication processes are usually after the exact deduplication and approximate deduplication processes, removing the semantic duplicates by clustering the embedding points and keeping representative data in each clusters. However, how to choose an appropriate pre-trained foundation models and whether this models can embed semantic information efficiently are still open questions.
@@
For Video multimodal data, video fingerprinting technology is uniquely applied to videos to identify and remove exact duplicates, addressing the challenge of video re-uploads and mirrored content, which is more prevalent and complex in videos than in static images(Xu et al., 2023c).
@@ 3.2.3. Data Enhancement
For vision-language models, enhancing the quality of image-caption datasets is crucial for training MLLMs. Improving the captions not only enhances their alignment with the images but also prevents the discarding of highly informative images due to poor text quality. Rewriting text using BLIP2 has been shown to effectively improve training results (Nguyen et al., 2023). However, for large-scale image-text datasets (approximately 1.28 billion), this method experiences diminishing returns in terms of ImageNet accuracy, possibly due to the lack of diversity in the generated captions (Nguyen et al., 2023). Approaches like LaCLIP (Fan et al., 2024) and VeCLIP (Lai et al., 2023a) employ LLMs with carefully designed prompts to generate more diverse captions. ShareGPT4V (Chen et al., 2023e) employed the 100K high-quality captions generated by GPT4-Vision to fine-tune an alternative caption model and named it as Share-Captioner. The Share-Captioner is capable of generating highly content-related descriptions with unified instruction for the pre-train dataset. MLM (Wang et al., 2024d) used GPT-4 or GPT-4V to constructing multimodal instruction tuning data on proposed quality scoring tasks to fine-tune MLM to realize accurate quality assessment. Then they adopt the fine-tuned MLM Filter to generate quality scores for each data point in the data pool and then select the high-quality data.
Improving image resolution can enhance the performance of MLLMs. Initially, MLLMs primarily processed fixed, lower-resolution inputs, typically around 224 pixels (Liu et al., 2024b; Chen et al., 2023j; Zhu et al., 2023). Recent models such as LLaVA-1.5 (Liu et al., 2023c) and BLiVA (Hu et al., 2024) have improved performance by increasing the input resolution to 336 pixels and integrating task-specific global features. Furthermore, models like Qwen-VL (Bai et al., 2023a) and OtterHD (Li et al., 2023n) have pushed resolution support to 448 pixels, incorporating fine-tuning in the visual encoder during training, while maintaining the original image size during inference, leading to more precise segmentation recognition. Notably, Monkey (Li et al., 2023k) has significantly raised the resolution to 896 pixels by employing multiple visual encoders and leveraging fine-tuning techniques from Qwen-VL. Models like LLAVA HR (Luo et al., 2024) and Vary (Wei et al., 2023) have introduced additional visual encoders to capture more complex features, requiring extensive pre-training tasks. LLAVA UHD (Xu et al., 2024) and Ureader (Ye et al., 2023a) have improved the model’s capacity to understand detailed image features through adaptive region segmentation. However, a fundamental challenge remains within the MLLM architecture: models using lower-resolution inputs struggle to detect fine details, whereas those with higher resolutions may underperform in tasks requiring a broader global understanding.
@@ 3.3. Commonly-used Datasets for MLLMs
This section we will briefly introduce commonly used multimodal datasets from different modality. The comprehensive summary of text-only datasets has been elaborately discussed in previous literature (Liu et al., 2024a); therefore, this part will not be extensively covered within the scope of this discussion. For the details of datasets introduced in this section, please refer to A.1.
@@ 3.3.1. Image Datasets
For vision langauge models, the image-text datasets commonly used can be categorized into three distinct types: image-caption datasets, interleaved image-text datasets, and visual question answering datasets.
@@
Image-caption pairs datasets are the most commonly used datasets in the pre-training stage of MLLMs. By leveraging the captions of each image, one can align the image modality with the text modality, enabling LLMs to understand image information. General image-caption datasets contain images with short captions, typically one sentence or a few words, that describe the key features of the image. These datasets vary significantly in size and scope, making them adaptable for various research and application needs. They are vital for the training and research of multimodal language-vision models, as they provide a corpus of image-text pairings that facilitate the understanding and generation of textual descriptions from visual inputs. Hence, collecting high-quality image-caption pairs is crucial in this phase. Fortunately, previous research in computer vision has amassed plenty of high-quality image-caption pairs.
@@
Notably, the LAION series (comprising LAION-400M (Schuhmann et al., 2021) and LAION-5B (Schuhmann et al., 2022)) represents some of the largest compilations, aggregating billions of pairs from the CommonCrawl archive. These datasets emphasize the maintenance of high quality through meticulous filtering based on the relevance of image-text pairs and the exclusion of inappropriate content. Derived from the LAION-5B subset, the LAION-COCO (Schuhmann et al., 2022) dataset is specifically designed to explore the impact of synthetic captions on model training. Also curated from CommonCrawl project, the COYO-700M (Minwoo Byeon, 2022) dataset emphasizes the informative correlation between images and alt-texts. The DataComp (Gadre et al., 2024) challenge introduces COMMONPOOL, a substantial multimodal dataset constructed from 12.8 billion image-text pairs that encourages innovation in dataset design.
Self-crawled captioning datasets including SBU (Ordonez et al., 2011), CC3M (Sharma et al., 2018), CC12M (Changpinyo et al., 2021), Wukong (Gu et al., 2022), ALT200M (Hu et al., 2022) and LTIP (Alayrac et al., 2022). Among these datasets, SBU (Ordonez et al., 2011) is the smallest containing only 1 million image-caption pairs curated from Flickr. Several datasets are curated from images with alt attributes from webpages. For example, Conceptual Captions (CC3M (Sharma et al., 2018) and CC12M (Changpinyo et al., 2021)) use generalized alt-text from the web and are processed with multiple filters, containing 3.3 million and 12 million image-text pairs respectively. Also leveraging alt-text for images on webpages, ALT200M (Hu et al., 2022) dataset collects 200 million images with their alt attributes, aiming for understanding scaling of vision-language models. For multilingual image-caption datasets, a representative one is the Wukong (Gu et al., 2022) dataset, which addresses the scarcity of large-scale datasets in the Chinese language, offering 100 million quality-assured pairs that are instrumental for the development of Chinese vision-language pre-training models. Larger dataset such as Long text & image pairs (LTIP) (Alayrac et al., 2022) dataset provides a unique collection of 312 million images with lengthier textual descriptions, enhancing models like Flamingo that are tailored for complex multimodal tasks. Collectively, these datasets advance the understanding of the complex interactions between visual and textual data, supporting the development of sophisticated artificial intelligence models.
Content Descriptive Image-Caption Dataset
Content descriptive image-caption datasets contain longer caption with more descriptive information. Usually each image in content description image-caption datasets contain at least 5 sentences of captions. Content descriptive image-caption datasets are essential tools for advancing the interface between visual content and linguistic descriptions, enhancing the ability of MLLMs to interpret and generate nuanced narratives from images. These datasets vary considerably in scope and methodology but share the goal of improving the semantic association between images and accompanying text. For example, MS-COCO (Chen et al., 2015) employs images with multiple reference captions to refine evaluation metrics, while Flickr30K (Young et al., 2014) extends the variety of descriptions to bolster semantic inference capabilities.
In addition, datasets like Visual Genome (Krishna et al., 2017) provide comprehensive annotations beyond simple captions, including objects, attributes, and relationships, to facilitate complex scene understanding. On the other hand, the AI Challenger Captions (Wu et al., 2017) dataset specifically addresses the need for non-English language representation in image captioning, offering extensive annotations in Chinese to bridge the semantic gap between low-level visual features and high-level conceptual descriptions.
Narrative and textual comprehension are further introduced by datasets like VIST (Huang et al., 2016), which emphasizes storytelling through visual sequences. TextCaps (Sidorov et al., 2020) incorporates reading comprehension into captioning. Collectively, these datasets not only propel advancements in automated image captioning but also tackle broader challenges in AI’s capability to process and generate meaningful visual and textual data, significantly enhancing context-aware machine understanding.
@@
A fundamental contrast between image-text caption datasets and interleaved image-text document datasets lies in their composition: image-caption pair datasets typically consist of a single image accompanied by multiple closely related captions, whereas interleaved image-text datasets consist of a text document interspersed with several illustrative images, wherein the correlation between the images and the accompanying text tends to be relatively lower. Interleaved image-text document datasets commonly used include M3W (Alayrac et al., 2022), MMC4 (Zhu et al., 2024), and OBELICS (Laurençon et al., 2024), which play crucial roles in advancing MLLMs (McKinzie et al., 2024). M3W (Alayrac et al., 2022) leverages data from 43 million web pages, integrating images with text to train models in few-shot learning scenarios. MMC4 (Zhu et al., 2024), an extension of the C4 corpus, aligns images with texts using CLIP features, resulting in a dataset containing over 101.2 million documents aimed at enhancing vision and language models within context. OBELICS (Laurençon et al., 2024), sourced from Common Crawl, comprises 141 million web pages and concentrates on high-quality multimodal content, filtering data to improve model training in comprehending and generating complex multimodal interactions.
Visual Question Answer (VQA) Image-Caption Datasets
@@
Visual question answer (VQA) image-caption datasets facilitate advanced research in MLLMs by allowing models to interpret images and answer related questions. These datasets typically consist of images paired with corresponding questions and answers, aiming to develop and evaluate models’ ability to understand visual content and provide accurate responses. For example, VQAv2.0 (Goyal et al., 2019) expands on its forerunner by introducing complementary images that prompt different answers to the same question, effectively addressing biases and increasing the dataset’s diversity. Visual-7W (Zhu et al., 2016) extends the scope by incorporating questions across multiple dimensions—what, where, when, who, why, how, and which—linked to specific objects within images. ST-VQA (Biten et al., 2019) integrates scene text into the VQA framework, making it possible to answer text-based questions from visual data. Shikra-RD (Chen et al., 2023j) leverages advanced language models to annotate images with relational descriptions, enhancing image comprehension. OCR-VQA focuses on reading text from book covers to answer related questions, thus combining elements of optical character recognition and VQA. DocVQA (Mathew et al., 2021) targets document images for extractive question answering, emphasizing precision in answers derived from visible text. A-OKVQA (Marino et al., 2019) introduces questions that demand commonsense and world knowledge, pushing the envelope on reasoning capabilities required from AI systems. TextVQA (Singh et al., 2019) specifically challenges models to read and understand text within images to respond accurately, and GQA (Hudson and Manning, 2019) promotes advanced reasoning over detailed visual scenes annotated with comprehensive scene graphs. These datasets collectively advance the field by enabling more nuanced interactions between AI models and the rich content within images, aiming for greater depth in the understanding and contextual integration of visual and textual data.
@@ 3.3.2. Video Datasets
Video datasets are comprehensive collections of video clips accompanied by associated annotations or labels. These datasets are meticulously designed to facilitate the training and evaluation of models for a wide range of video-related tasks, including video classification, object detection, video question answering, action detection and video captioning.
In the realm of video datasets, MSR-VTT (Xu et al., 2016), WebVid-2M, WebVid-10M (Bain et al., 2021), and VTP (Alayrac et al., 2022) are three prominent examples commonly utilized for pre-training MLLMs. Specifically, the MSR-VTT (Bain et al., 2021) dataset features 10,000 video clips totaling 41.2 hours, each annotated with multiple descriptive sentences. Similarly, the WebVid series offers millions of web-sourced video clips with accompanying captions, expanding from 2.5 million pairs in WebVid-2M to 10 million pairs in WebVid-10M (Bain et al., 2021), covering a vast array of 13,000 hours of video content. The VTP (Alayrac et al., 2022) dataset further enriches the field with 27 million short video-text pairs sourced from a select few high-quality websites. Moreover, the recently introduced Panda-70M (Chen et al., 2024c) dataset contains a massive 70 million high-quality video-caption pairs, designed to enhance the training of high-performance MLLMs. Collectively, these datasets are integral to advancing the capabilities of MLLMs in understanding and interacting with video content.

### v2

@@ 3.2.2. Deduplications
Existing deduplication methods consider exact duplication, approximate duplication and semantic duplication in sentence-level (or sequence-level) and document-level. Exact deduplication is the most simple way to remove duplication. For sentence or sequence level, exact deduplication consider to remove exact string matching between sentence (or sequence) (Raffel et al., 2020; Suárez et al., 2019). To improve computational and memory efficiency, Suffix Arrays (Manber and Myers, 1993; Lee et al., 2022) and Bloom Filters (Computer, 2023; Soldaini et al., 2024) are used to achieve parallelized linear time complexity. For document level exact deduplication, URL deduplication is considered to remove exactly same web pages (Penedo et al., 2023). Approximate deduplication methods mainly focus on document-level duplicates. Usually, Locality Sensitive Hashing (LSH)-based MinHash (Broder, 1997) or SimHash (Charikar, 2002) methods are used to remove approximately duplicate documents (Computer, 2023; Lee et al., 2022; Penedo et al., 2023). These methods can achieve document-level deduplication with linear time and space complexity of the document number and can be implemented in a highly distributed setting (Lee et al., 2022). Apart from hashing-based approximate deduplication, recently there are some work consider to leverage pre-trained foundation models as semantic embedding metric for document-level embedding (Silcock et al., 2022; Kaddour, 2023; Abbas et al., 2023; Tirumala et al., 2023). These Semantic deduplication methods consider using Sentence-BERT (Reimers and Gurevych, 2019) MPNET (Song et al., 2020) for embedding (Silcock et al., 2022), E5-Large (Wang et al., 2022) for embedding (Kaddour, 2023), and OPT-125M (Zhang et al., 2022) for embedding (Abbas et al., 2023; Tirumala et al., 2023). These semantic deduplication processes are usually after the exact deduplication and approximate deduplication processes, removing the semantic duplicates by clustering the embedding points and keeping representative data in each clusters. However, how to choose an appropriate pre-trained foundation models and whether this models can embed semantic information efficiently are still open questions.
@@
For Video multimodal data, video fingerprinting technology is uniquely applied to videos to identify and remove exact duplicates, addressing the challenge of video reuploads and mirrored content, which is more prevalent and complex in videos than in static images(Xu et al., 2023c).
@@ 3.2.3. Data Enhancement
For vision-language models, enhancing the quality of image-caption datasets is crucial for training MLLMs. Improving the captions not only enhances their alignment with the images but also prevents the discarding of highly informative images due to poor text quality. Rewriting text using BLIP2 has been shown to effectively improve training results (Nguyen et al., 2023). However, for large-scale image-text datasets (approximately 1.28 billion), this method experiences diminishing returns in terms of ImageNet accuracy, possibly due to the lack of diversity in the generated captions (Nguyen et al., 2023). Approaches like LaCLIP (Fan et al., 2024) and VeCLIP (Lai et al., 2023a) employ LLMs with carefully designed prompts to generate more diverse captions. ShareGPT4V (Chen et al., 2023e) employed the 100K high-quality captions generated by GPT4-Vision to fine-tune an alternative caption model and named it as Share-Captioner. The Share-Captioner is capable of generating highly content-related descriptions with unified instruction for the pre-train dataset. MLM (Wang et al., 2024c) used GPT-4 or GPT-4V to constructing multimodal instruction tuning data on proposed quality scoring tasks to fine-tune MLM to realize accurate quality assessment. Then they adopt the fine-tuned MLM Filter to generate quality scores for each data point in the data pool and then select the high-quality data.
Improving image resolution can enhance the performance of MLLMs. Initially, MLLMs primarily processed fixed, lower-resolution inputs, typically around 224 pixels (Liu et al., 2024b; Chen et al., 2023j; Zhu et al., 2023). Recent models such as LLaVA-1.5 (Liu et al., 2023c) and BLiVA (Hu et al., 2024) have improved performance by increasing the input resolution to 336 pixels and integrating task-specific global features. Furthermore, models like Qwen-VL (Bai et al., 2023a) and OtterHD (Li et al., 2023m) have pushed resolution support to 448 pixels, incorporating fine-tuning in the visual encoder during training, while maintaining the original image size during inference, leading to more precise segmentation recognition. Notably, Monkey (Li et al., 2023j) has significantly raised the resolution to 896 pixels by employing multiple visual encoders and leveraging fine-tuning techniques from Qwen-VL. Models like LLAVA HR (Luo et al., 2024) and Vary (Wei et al., 2023) have introduced additional visual encoders to capture more complex features, requiring extensive pre-training tasks. LLAVA UHD (Xu et al., 2024) and Ureader (Ye et al., 2023a) have improved the model’s capacity to understand detailed image features through adaptive region segmentation. However, a fundamental challenge remains within the MLLM architecture: models using lower-resolution inputs struggle to detect fine details, whereas those with higher resolutions may underperform in tasks requiring a broader global understanding.
@@ 3.3. Commonly-used Datasets for MLLMs
In this section we will briefly introduce commonly used multimodal datasets from different modality. These datasets are not utilized for this research. The comprehensive summary of text-only datasets has been elaborately discussed in previous literature (Liu et al., 2024a); therefore, this part will not be extensively covered within the scope of this discussion.
@@ 3.3.1. Image Datasets
For vision langauge models, the commonly used image-text datasets can be categorized into three distinct types: image-caption datasets, interleaved image-text datasets, and visual question answering datasets.
@@
Image-caption pairs datasets are the most commonly used datasets in the pre-training stage of MLLMs. By leveraging the captions of each image, one can align the image modality with the text modality, enabling LLMs to understand image information.
@@
General image-caption datasets contain images with short captions, typically one sentence or a few words, that describe the key features of the image. Notably, the LAION series (comprising LAION-400M (Schuhmann et al., 2021) and LAION-5B (Schuhmann et al., 2022)) represents some of the largest compilations, aggregating billions of pairs from the CommonCrawl. These datasets emphasize the maintenance of high quality through meticulous filtering based on the relevance of image-text pairs and the exclusion of inappropriate content. Derived from the LAION-5B subset, the LAION-COCO (Schuhmann et al., 2022) dataset is specifically designed to explore the impact of synthetic captions on model training. Also curated from CommonCrawl project, the COYO-700M (Minwoo Byeon, 2022) dataset emphasizes the informative correlation between images and alt-texts. The DataComp (Gadre et al., 2024) challenge introduces COMMONPOOL, a substantial multimodal dataset constructed from 12.8 billion image-text pairs that encourages innovation in dataset design. Apart from CommonCrawl project, some self-crawled datasets also play a vital role. Several datasets are curated from images with alt attributes from webpages. For example, Conceptual Captions (CC3M (Sharma et al., 2018) and CC12M (Changpinyo et al., 2021)) use generalized alt-text from the web and are processed with multiple filters, containing 3.3 million and 12 million image-text pairs respectively. Also leveraging alt-text for images on webpages, ALT200M (Hu et al., 2022) dataset collects 200 million images with their alt attributes, aiming for understanding scaling of vision-language models. For multilingual image-caption datasets, a representative one is the Wukong (Gu et al., 2022) dataset, which addresses the scarcity of large-scale datasets in the Chinese language, offering 100 million quality-assured pairs that are instrumental for the development of Chinese vision-language pre-training models. Larger dataset such as Long text & image pairs (LTIP) (Alayrac et al., 2022) dataset provides a unique collection of 312 million images with lengthier textual descriptions, enhancing models like Flamingo that are tailored for complex multimodal tasks.
Content descriptive image-caption datasets contain longer caption with more descriptive information. Usually each image in content description image-caption datasets contain at least 5 sentences of captions. Content descriptive image-caption datasets are essential for advancing the interface between visual content and text descriptions, enhancing the ability of MLLMs to interpret and generate more detailed description from images. For example, MS-COCO (Chen et al., 2015) employs images with multiple reference captions to refine evaluation metrics, while Flickr30K (Young et al., 2014) extends the variety of descriptions to bolster semantic inference capabilities. In addition, datasets like Visual Genome (Krishna et al., 2017) provide comprehensive annotations beyond simple captions, including objects, attributes, and relationships, to facilitate complex scene understanding. While the AI Challenger Captions (Wu et al., 2017) dataset specifically addresses the need for non-English language representation in image captioning, offering extensive annotations in Chinese to bridge the semantic gap between low-level visual features and high-level conceptual descriptions. Narrative and textual comprehension are further introduced by datasets like VIST (Huang et al., 2016), which emphasizes storytelling through visual sequences. TextCaps (Sidorov et al., 2020) incorporates reading comprehension into captioning. Collectively, these datasets not only propel advancements in automated image captioning but also tackle broader challenges in AI’s capability to process and generate meaningful visual and textual data, significantly enhancing context-aware machine understanding.
@@
A fundamental contrast between image-text caption datasets and interleaved image-text document datasets lies in their composition: image-caption pair datasets typically consist of a single image accompanied by multiple closely related captions, whereas interleaved image-text datasets consist of a text document interspersed with several illustrative images. The correlation between the images and the accompanying text tends to be relatively lower, but these datasets usually provides more semantic information, which is crucial for maintaining models’ semantic understanding. M3W (Alayrac et al., 2022) leverages data from 43 million web pages, integrating images with text to train models in few-shot learning scenarios. MMC4 (Zhu et al., 2024), an extension of the C4 corpus, aligns images with texts using CLIP features, resulting in a dataset containing over 101.2 million documents. OBELICS (Laurençon et al., 2024), sourced from CommonCrawl, comprises 141 million web pages. OmniCorpus (Li et al., 2024) scales the interleaved image-text datasets to 10 billion-level, providing much richer image and text information from diverse sources.
Visual Question Answer (VQA) Datasets
@@
Visual question answer (VQA) datasets facilitate advanced research in MLLMs by allowing models to interpret images and answer related questions. These datasets typically consist of images paired with corresponding questions and answers, aiming to develop and evaluate models’ ability to understand visual content and provide accurate responses. For example, VQAv2.0 (Goyal et al., 2019) expands on its forerunner by introducing complementary images that prompt different answers to the same question, effectively addressing biases and increasing the dataset’s diversity. Visual-7W (Zhu et al., 2016) extends the scope by incorporating questions across multiple dimensions—what, where, when, who, why, how, and which—linked to specific objects within images. ST-VQA (Biten et al., 2019) integrates scene text into the VQA framework, making it possible to answer text-based questions from visual data. Shikra-RD (Chen et al., 2023j) leverages advanced language models to annotate images with relational descriptions, enhancing image comprehension. OCR-VQA focuses on reading text from book covers to answer related questions, combining elements of optical character recognition and VQA. DocVQA (Mathew et al., 2021) targets document images for extractive question answering, emphasizing precision in answers derived from visible text. A-OKVQA (Marino et al., 2019) introduces questions that demand commonsense and world knowledge, pushing the envelope on reasoning capabilities required from AI systems. TextVQA (Singh et al., 2019) specifically challenges models to read and understand text within images to respond accurately, and GQA (Hudson and Manning, 2019) promotes advanced reasoning over detailed visual scenes annotated with comprehensive scene graphs. These datasets collectively advance the field by enabling more nuanced interactions between AI models and the rich content within images, aiming for greater depth in the understanding and contextual integration of visual and textual data.
@@ 3.3.2. Video Datasets
Video datasets are comprehensive collections of video clips accompanied by associated annotations or labels. These datasets are meticulously designed to facilitate the training and evaluation of models for a wide range of video-related tasks, including action recognition, various video-text tasks, as well as video-centric dialogue.
Commonly used video-text dataset such as the MSR-VTT (Bain et al., 2021) features 10,000 video clips totaling 41.2 hours, each annotated with multiple descriptive sentences. Similarly, the WebVid series offers millions of web-sourced video clips with accompanying captions, expanding from 2.5 million pairs in WebVid-2M to 10 million pairs in WebVid-10M (Bain et al., 2021), covering a vast array of 13,000 hours of video content. The VTP (Alayrac et al., 2022) dataset further enriches the field with 27 million short video-text pairs sourced from a select few high-quality websites. Moreover, the recently introduced Panda-70M (Chen et al., 2024c) dataset contains a massive 70 million high-quality video-caption pairs, designed to enhance the training of high-performance MLLMs. Collectively, these datasets are integral to advancing the capabilities of MLLMs in understanding and interacting with video content. InternVid (Wang et al., 2023a) scales video-text dataset to 230M annotated video-text pairs, with every video lasting 351.9s on average.